### v1

@@ 3.1 Bias Analysis
$\mathbb{E}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-g(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})\bigg{]}=\mathbb{E}[f(\bm{Y})]$
Proof 3.2
[…]
@@ 3.2 Variance Analysis
Definition 3.3
Given function $f$ that takes input $\bm{Y}$ and $g$ that takes input $\bm{X}$ (so both $f,g$ are considered fixed here), we denote
$\sigma_{f-g}^{2}:=\mathrm{Var}(f(\bm{Y})-g(\bm{X})),\text{ }\sigma^{2}_{f}:=\mathrm{Var}(f(\bm{Y})),\text{ and }\sigma^{2}_{g}:=\mathrm{Var}(g(\bm{X})).$
Similarly, we denote cost of generating a coupled sample $f(\bm{Y})-g(\bm{X})$ as $c_{f-g}$, the cost of generating $f(\bm{Y})$ as $c_{f}$ and the cost of generating $g(\bm{X})$ as $c_{g}$.
Lemma 3.4
Using the notations from 3.3, we have
$\mathrm{Cov}(f(\bm{Y}),g(\bm{X}))=\frac{\sigma^{2}_{f}+\sigma^{2}_{g}-\sigma^{2}_{f-g}}{2}.$
We make two remarks here. First, since we generally obtain $\bm{X}$ in the same process during the generation of $\bm{Y}$, thus typically in either sense of the cost, we have $c_{f-g}\approx c_{f}$. Second, the term cost is intentionally left ambiguous because its interpretation varies depending on the context. In some cases, cost refers to the sample size, while in others it represents the wall-clock time required to generate and evaluate the sample. Furthermore, in certain scenarios, both time and computational resources must be jointly considered and optimized, under parallelization and other speedup, making the concept of cost more instance dependent. Therefore, for now, we use cost as an umbrella term to capture various meanings under different contexts.
Lemma 3.5
@@
Proof 3.6
[…]
Having established the unbiasedness and the variance form of the PEMC estimator, we could turn to inference. In particular, using consistent estimates of variance, the application of the central limit theorem enables the construction of valid asymptotic confidence intervals. Moreover, a range of inferential techniques may be employed—such as deriving high-probability bounds or other non-asymptotic guarantees—the asymptotic confidence interval construction suffices to illustrate the point. The approach follows closely from results in prediction-powered inference Angelopoulos et al. (2023).
Theorem 3.7 (Asymptotic Confidence Intervals)
In the set up of the evaluation algorithm 2, given the prediction model $g$, $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$ and $(\bm{\tilde{X}}_{j})_{j\in[N]}$, we define
$\hat{\sigma}^{2}_{f-g}=\frac{1}{n}\sum_{i=1}^{n}\Bigl{(}f(\bm{Y}_{i})-g(\bm{X}_{i})-\Bigl{(}\frac{1}{n}\sum_{i^{\prime}=1}^{n}f(\bm{Y}_{i^{\prime}})-g(\bm{X}_{i^{\prime}})\Bigr{)}\Bigr{)}^{2},$
@@
$\hat{\sigma}^{2}_{g}=\frac{1}{N}\sum_{j=1}^{N}\bigl{(}g(\bm{\tilde{X}}_{j})-\frac{1}{N}\sum_{j^{\prime}=1}^{N}g(\bm{\tilde{X}}_{j^{\prime}})\bigr{)}^{2},$
@@
is an asymptotically valid $1-\alpha$ confidence interval for $\mathbb{E}[f(\bm{Y})]$, as $n,N\to\infty$.
Proof 3.8
[…]
3.3 Variance Reduction
@@
[…]
In practice, $c_{f-g}\gg c_{g}$ often holds which suggests one should set $N\gg n$ according to Lemma 3.9. While Lemma 3.9 relies on treating $n,N\in\mathbb{R}^{+}$ as continuous—ignoring the integer constraints $n,N\in\mathbb{N}^{+}$— this continuous approximation still provides valuable guidance when $n,N\geq 1$. In fact, in practice, we found the ratio suggested by Lemma 3.9 guides near-optimal parameter selection. However, the standard MC corresponds to the case $N=0$, a scenario not covered by Lemma 3.9, and thus requires separate consideration. Building on Lemma 3.9, we can estimate an upper bound on the variance reduction PEMC can achieve relative to standard MC under these idealized assumptions.
Lemma 3.11
Assume $c_{f-g}=c_{f}$. In the same setup as Lemma 3.9, the variance ratio between PEMC under the optimal allocation and standard MC follows as
@@
Proof 3.12
[…]
At this point, a natural question arises: how can we gauge $\frac{\sigma_{f-g}^{2}}{\sigma_{f}^{2}}$ or $\frac{\sigma_{g}^{2}}{\sigma_{f}^{2}}$ in practice? How are these ratios tied to the quality of the predictive model $g$ in PEMC? As we shall see, the extent of variance reduction that PEMC can deliver is linked to how well $g$ is trained. To illustrate this, we first consider an ideal scenario where $g$ is trained to the optimum, i.e., the true regression function.
Lemma 3.13
Suppose $c_{f-g}=c_{f}$, $f(\bm{Y})$ is square-integrable and $g=\mathbb{E}[f(\bm{Y})\,|\,\bm{X}]$. Define $\rho:=\mathrm{corr}(f,g)$ and $c:=\frac{c_{g}}{c_{f}}$. Then we have $\rho=\frac{\sigma_{g}}{\sigma_{f}}$ and the $\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}$ in Lemma 3.11 becomes
@@
Proof 3.14
Proof When $g=\mathbb{E}[f(\bm{Y})\,|\,\bm{X}]$ is the true regressor, i.e.,
[…]
@@
Lemma 3.13 says that the variance reduction of PEMC relative to standard MC can be approximated by the function $r(\rho,c)$ in (8), where $\rho=\sigma_{g}/\sigma_{f}$ captures the predictive power of $g(\bm{X})$ for $f(\bm{Y})$, and $c=c_{g}/c_{f}$ quantifies the relative cost of evaluating $g(\bm{X})$ versus generating full samples for $f(\bm{Y})$. In Figure 3.3, we visualize the variance reduction function $r(\rho,c)$ for $\rho,c\in(0,1)$. On the left is a contour plot of $r(\rho,c)$ where the red line traces the curve $r(\rho,c)=1$, corresponding to the “break-even” boundary where PEMC’s variance matches that of MC. Regions below this line, $\{(\rho,r)\in[0,1]^{2}|r(\rho,c)<1\}$, represent regimes of $\rho$ and $c$ where PEMC achieves variance reduction, i.e., where $\rho$ is sufficiently large and $c$ is sufficiently small. On the right, the graph of $r(\rho,c)$ as a function of $\rho$ when $c=0.001$ provides a clear benchmark: a correlation $\rho=0.5$ yields approximately 22.2% variance reduction, while $\rho=0.7$ yields about 45.8%. In our PEMC applications, we design feature $\bm{X}$ and predictive model $g$ so that, for most $\bm{\theta}\in\Theta$ during evaluation, we can safely gauge that $c$ falls between $10^{-2}$ to $10^{-3}$ and $\rho$ exceeds 0.5. In these regions, Lemma 3.9 suggests choosing $N/n$ in the range of 5 to 20 for near-optimal performance. In our experiments, a ratio of $N=10n$ proved effective.
[…]
@@
Variance Reduction Function $r(\rho,c)$. Left figure: A contour map of the variance reduction ratio $r(\rho,c)$ as a function of the correlation $\rho=\sigma_{g}/\sigma_{f}$ and relative cost $c=c_{g}/c_{f}$ The red curve indicates the curve $r(\rho,c)=1$ and the regimes where PEMC attains variance reduction over MC lie below the line. Right figure: A graph of $r(\rho,c)$ as a function of $\rho$ when $c=10^{-3}$.
3.3.1 Learning Theory Estimates
The statements in Lemma 3.9 assume that $g$ represents the true regression function $\mathbb{E}[f|\bm{X}]$. In practice, however, $g$ is obtained through a learning procedure applied to finite training samples, and thus will only approximate the true conditional expectation. Consequently, any practical implementation of PEMC must account for the approximation and estimation errors inherent in the training process. Moreover, as the parameter $\bm{\theta}$ varies, the correlation $\rho$ between $g(\bm{X})$ and $f(\bm{Y})$, the complexity of the underlying model and payoff function also changes.
@@ 3.3.1 Learning Theory Estimates
To rigorously quantify these effects, one can invoke tools from statistical learning theory, such as uniform convergence bounds, VC-dimension, or Rademacher complexities (Vapnik 1998, Bartlett and Mendelson 2005, Mohri et al. 2018). These quantities, by relating the complexity of the function class used to represent $g$ and the available training sample size $N_{\text{train}}$, lead to error bounds that ensure, with high probability, that the trained predictor $g$ is close to the true regression function within a controlled margin. Although numerous, more refined approaches exist, we avoid delving deeply into them so as not to distract from our main theme. Instead, we adopt a standard, off-the-shelf, almost ‘placeholder’-type result detailed in the Appendix, that illustrates the typical form of guarantees one could expect from PEMC.
Lemma 3.15
@@
The simplified version captures the essential idea: under suitable regularity conditions, one can achieve the result of Lemma 3.9 within an arbitrarily small margin, as both the sample size $N_{\text{train}}$ and the complexity of the hypothesis class increase.
3.4 Control Variate Coefficient
@@
$PEMC(a):=\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-ag(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}ag(\bm{\tilde{X}}_{j})$ (9)
@@
The introduction of $a$ is analogous to CV approaches, where the parameter $a$ is chosen to minimize variance. In PEMC, this viewpoint may appear somewhat redundant since $ag$ is also a valid prediction model, which suggests $a$ would be chosen implicitly during the training of the prediction model. Nevertheless, suppose $g$ is fixed. Then $\frac{1}{n}\sum_{i=1}^{n}-g(\bm{X}_{i})+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})$ can be viewed as a readily available, zero-mean control variate. We can then choose $a$ which minimizes the variance. The optimal $a^{*}$ then is:
$a^{*}=\frac{\mathrm{Cov}(f(\bm{Y}),g(\bm{X}))}{(n/N+1)\mathrm{Var}(g(\bm{X}))}.$
@@
In the ideal scenario, we have $N\gg n$ and $g(\bm{X})=\mathbb{E}[f(\bm{Y})|\bm{X}]$, which would imply $\mathrm{Cov}(f(\bm{Y})-g(\bm{X}),g(\bm{X}))=0$ and $\mathrm{Cov}(f(\bm{Y}),g(\bm{X}))=\mathrm{Var}(g(\bm{X}))$. This would further imply $a^{*}=\frac{1}{1+n/N}\approx 1$. Consequently, we can justify PEMC’s default choice of $a=1$, which saves us the trouble of estimating $a^{*}$, and gives PEMC a degree of robustness.
3.5 Variance Reduction of PEMC: An Example on Asian Options
We revisit the Asian option example and conduct experiments using the arithmetic Asian call option under the Geometric Brownian Motion (GBM) model. We choose the GBM model instead of the Heston model (1) to facilitate a direct comparison between PEMC, standard Monte Carlo (MC), and the traditional Control Variate (CV) method on the extent of variance reduction. This is because the geometric Asian option, which serves as a well-established control variate, is analytically tractable within the GBM framework but lacks a closed-form solution under the Heston model (Dingec et al. 2015).
This experiment aims to demonstrate the typical level of variance reduction that PEMC achieves. Our focus is strictly limited to variance reduction, because the GBM model actually allows for highly efficient and parallelized generation of both asset paths and corresponding payoffs, making the simulation of $\bm{Y}$ computationally straightforward. In other words, one could not fully leverage PEMC’s strengths in GBM model and this toy experiment only serves to provide insights into PEMC’s effectiveness in a controlled setting where a known CV exists. Practical applications and actual use-cases of PEMC will be provided in Section 4.
3.5.1 Experimental Setup
$dS_{t}=$ $rS_{t}dt+\sigma S_{t}dW_{t}^{S}.$ (10)
The payoff calculation $P_{A}(\{S_{t}\}_{t})=(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0)^{+}$ with $n_{D}=252$, where simulations are carried out using daily time increments (i.e., $\Delta t=1$). The parameter space $\bm{\theta}:=(r,S_{0},\sigma,K)\in\bm{\Theta}$ is: $r\in[0.01,0.03]$, $S_{0}\in[80,120]$, $\sigma\in[0.05,0.25]$, and $K\in[90,110]$. For our feature $\bm{X}$, we experiment with the whole sum of Brownian increments $W_{T}^{S}:=\sum_{j=1}^{252}\Delta W^{S}_{j}$ with $\dim(\bm{X})=1$. We also tried a more granular representation where we partition 252 Brownian increments evenly into 14 parts, each summing 18 consecutive increments: $[\bm{X}]_{i}=\sum_{j=18(i-1)+1}^{18i}\Delta W^{S}_{j}$ for $i\in[14]$, yielding $\dim(\bm{X})=14$. The neural network is trained on a dataset comprising $N_{\text{train}}=1.28\times 10^{6}$ samples, generated by uniformly sampling $\bm{\theta}\sim\bm{\Theta}$.
The neural network architecture consists of three components: a $\bm{\theta}$ network branch, a $\bm{X}$ network branch, and a combined network. This architecture incorporates several modern deep learning practices, including Batch Normalization (Ioffe and Szegedy 2015), Rectified Linear Unit (ReLU) activations (Nair and Hinton 2010), and skip connections (He et al. 2016) to enhance training stability and model performance. The $\bm{\theta}$ network branch processes the 4-dimensional input $\bm{\theta}$ using two fully connected layers with Batch Normalization and ReLU activation functions, outputting a 10-dimensional feature vector for subsequent processing. The $\bm{X}$ network branch uses two fully connected layers with width $\max(32,2\dim(\bm{X}))$ neurons. Finally, the combined network integrates outputs from both branches through concatenation and employs skip connections featured in ResNet architectures to preserve information flow. The concatenated features pass through two additional fully connected layers with Batch Normalization and ReLU activation to produce the final prediction for PEMC. The network is implemented using PyTorch (Paszke et al. 2019) and trained using the Adam optimizer (Kingma and Ba 2014) with a learning rate of $1\times 10^{-3}$, incorporating dropout layers (Srivastava et al. 2014) with a rate of 0.5 after each hidden layer to prevent overfitting.
3.5.2 Evaluation
$P_{G}(\{S_{t}\}_{t})=((\prod_{i=1}^{n_{D}}S_{t_{i}})^{\frac{1}{{n_{D}}}}-K,0)^{+},$
for the evaluation at $\bm{\theta}=(r,S_{0},\sigma,K)=(0.02,100,0.2,100)$. The CV estimator is given by
$\text{CV}=\frac{1}{n}\sum_{i=1}^{n}\left(P_{A}(\{S_{t}^{(i)}\}_{t})-P_{G}(\{S_{t}^{(i)}\}_{t})\right)+P_{G}^{\text{exact}}(\bm{\theta}),$
[…]
Table 1: Root Mean Squared Error from 300 Experiments
Method $n=1000$ $n=4000$ $n=9000$
Monte Carlo (MC) 0.2376 0.1173 0.0854
PEMC ($\dim\bm{X}=1$) 0.1509 0.0809 0.0481
PEMC ($\dim\bm{X}=14$) 0.0781 0.0397 0.0261
Geometric CV 0.0099 0.0051 0.0036
[…]
Performance of Estimators for Asian Options. Comparison of estimator across Monte Carlo (MC), PEMC, and Geometric Control Variate (CV) based on 300 experiments. Top: Mean squared errors (MSE) plot as a function of $n$ for (left) PEMC with $\dim\bm{X}=1$ and (right) PEMC with $\dim\bm{X}=14$, both compared against MC and Geometric CV. Bottom: Corresponding boxplots of the 300 estimates.
Given that all estimators are unbiased, our analysis focuses exclusively on variance-related metrics. The experimental results demonstrate the superior performance of the PEMC framework compared to the standard Monte Carlo (MC) estimator. In particular, in all sample sizes, the PEMC framework delivers visible variance reduction over MC: PEMC’s basic variant with $\dim\bm{X}=1$ achieves a 30-40% reduction in root mean squared error (RMSE) relative to MC, while the more sophisticated variant with $\dim\bm{X}=14$ attains an impressive 65-70% reduction in RMSE over MC. However, the Geometric CV emerges as the most efficient estimator by far, with RMSE an order of magnitude smaller than MC.
To better understand this hierarchy in estimator performance, it is helpful to recognize PEMC not as a specific estimator, but rather as a framework that latches on and enhance an existing baseline. In our current example, we chose standard MC as the baseline, and design our PEMC directly upon MC to improve it. This also illustrates a common state of practice: in most complex scenarios, sophisticated baselines like CV are not available, and practitioners must rely on some variants of naive MC methods. Then, PEMC can be applied to these baselines to achieve variance reduction. However, PEMC is a flexible framework designed to enhance general simulation baselines, including sophisticated ones, not just standard Monte Carlo. In particular, when a control variate method is available, such as in our current example, PEMC can be built upon it to potentially achieve even greater variance reduction. To see how, in Algorithm 1, we change the label from $P_{A}(\{S_{t}\}_{t})$ (MC baseline) to $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\bm{\theta})$ (Geometric CV baseline), and let the model $g$ learn
$g(\bm{\theta},\bm{X}(\bm{\theta}))\approx$ $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}\bigg{[}P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\bm{\theta})\bigg{|}(\bm{\theta},\bm{X}(\bm{\theta}))\bigg{]}$
$=$ $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}\bigg{[}P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})\bigg{|}(\bm{\theta},\bm{X}(\bm{\theta}))\bigg{]}+P_{G}^{\text{exact}}(\bm{\theta}).$ (11)
This formulation suggests we can train $g$ on the difference of $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})$ as labels and then add back the closed-form $P_{G}^{\text{exact}}(\bm{\theta})$ to $g$. We implemented this approach, which we call “Boost PEMC,” using the basic PEMC variant where $\bm{X}=W_{T}^{S}:=\sum_{j=1}^{252}\Delta W^{S}_{j}$ with $\dim\bm{X}=1$. As shown in Table 2 and Figure 3.5.2, this Boost PEMC indeed achieves variance reduction over the sophisticated Geometric CV estimator, delivering a 35-40% reduction in RMSE, remarkably consistent with the relative performance gain we observed when applying PEMC to the MC baseline in Figure 3.5.2.
Table 2: Root Mean Squared Error from 300 Experiments
Method $n=1000$ $n=4000$ $n=9000$
Boost PEMC ($\dim\bm{X}=1$) 0.0065 0.0031 0.0021
Geometric CV 0.0099 0.0051 0.0036
[…]
Performance of CV-based Estimators for Asian Options. Comparison of estimator performance for Boost PEMC and Geometric Control Variate (CV) based on 300 experiments. Left: Mean squared errors (MSE) plot as a function of $n$. Right: Corresponding boxplots of 300 estimates.
4 Applications

### v2

@@ 4.1. Bias Analysis
$\mathbb{E}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}(f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\tilde{X}}_{j})\bigg{]}=\mathbb{E}[f(\boldsymbol{Y})]$
Proof.
Proof First note that $g$ is pre-trained. The proof then follows since $\boldsymbol{X}_{i}$ and $\boldsymbol{\tilde{X}}_{j}$ have the same marginal distribution for any $i,j$. ∎
@@ 4.2. Variance Analysis
Definition 2.
Given function $f$ that takes input $\boldsymbol{Y}$ and $g$ that takes input $\boldsymbol{X}$ (so both $f,g$ are considered fixed here), we denote
$\sigma_{f-g}^{2}:=\mathrm{Var}(f(\boldsymbol{Y})-g(\boldsymbol{X})),\text{ }\sigma^{2}_{f}:=\mathrm{Var}(f(\boldsymbol{Y})),\text{ and }\sigma^{2}_{g}:=\mathrm{Var}(g(\boldsymbol{X})).$
Similarly, we denote cost of generating a coupled sample $f(\boldsymbol{Y})-g(\boldsymbol{X})$ as $c_{f-g}$, the cost of generating $f(\boldsymbol{Y})$ as $c_{f}$ and the cost of generating $g(\boldsymbol{X})$ as $c_{g}$.
Lemma 1.
Using the notations from 2, we have
$\mathrm{Cov}(f(\boldsymbol{Y}),g(\boldsymbol{X}))=\frac{\sigma^{2}_{f}+\sigma^{2}_{g}-\sigma^{2}_{f-g}}{2}.$
We make two remarks here. First, since we generally obtain $\boldsymbol{X}$ in the same process during the generation of $\boldsymbol{Y}$, thus typically in either sense of the cost, we have $c_{f-g}\approx c_{f}$. Second, the term cost is intentionally left unspecified because its interpretation varies depending on the context. In some context, cost refers to the sample size, while in others it represents the wall-clock time required to generate and evaluate all the samples, making the concept of cost instance dependent. Therefore, for now, we use cost as an umbrella term to capture various meanings under different contexts.
Lemma 2.
@@
Proof.
Proof First note that $g$ is pre-trained. Moreover, in the evaluation phase, data $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$ are generated independent of $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$. ∎
Having established the unbiasedness and the variance form of the PEMC estimator, we could turn to inference. In particular, using consistent estimates of variance, the application of the central limit theorem enables the construction of valid asymptotic confidence intervals. Moreover, a range of inferential techniques may be employed—such as deriving high-probability bounds or other non-asymptotic guarantees—the asymptotic confidence interval construction suffices to illustrate the point. The approach follows closely from results in prediction-powered inference [6].
Theorem 3 (Asymptotic Confidence Intervals).
In the set up of the evaluation algorithm 2, given the prediction model $g$, $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$ and $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$, we define
$\hat{\sigma}^{2}_{f-g}=\frac{1}{n}\sum_{i=1}^{n}\Bigl{(}f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i})-\Bigl{(}\frac{1}{n}\sum_{i^{\prime}=1}^{n}f(\boldsymbol{Y}_{i^{\prime}})-g(\boldsymbol{X}_{i^{\prime}})\Bigr{)}\Bigr{)}^{2},$
@@
$\hat{\sigma}^{2}_{g}=\frac{1}{N}\sum_{j=1}^{N}\bigl{(}g(\boldsymbol{\tilde{X}}_{j})-\frac{1}{N}\sum_{j^{\prime}=1}^{N}g(\boldsymbol{\tilde{X}}_{j^{\prime}})\bigr{)}^{2},$
@@
is an asymptotically valid $1-\alpha$ confidence interval for $\mathbb{E}[f(\boldsymbol{Y})]$, as $n,N\to\infty$.
Proof.
Proof Theorem 1 establishes PEMC is unbiased. Moreover, $\hat{\sigma}^{2}_{f-g},\hat{\sigma}^{2}_{g}$ are consistent estimates of ${\sigma}^{2}_{f-g},{\sigma}^{2}_{g}$, the results follows from Lemma 2 and the central limit theorem and Slutsky’s theorem. ∎
4.3. Variance Reduction
@@
is jointly convex in $n,N\in\mathbb{R}^{+}$. The objective value also approaches infinity as $N\to 0$ or $n\to 0$. Consequently, the strict convexity ensures the existence of a global minimizer in the interior and directly solving the Lagrangian gives us the result. ∎
In practice, $c_{f-g}\gg c_{g}$ often holds which suggests one should set $N\gg n$ according to Lemma 3. While Lemma 3 relies on treating $n,N\in\mathbb{R}^{+}$ as continuous—ignoring the integer constraints $n,N\in\mathbb{N}^{+}$— this continuous approximation still provides valuable guidance when $n,N\geq 1$. In fact, in practice, we found the ratio suggested by Lemma 3 guides near-optimal parameter selection. However, the standard MC corresponds to the case $N=0$, a scenario not covered by Lemma 3, and thus requires separate consideration. Building on Lemma 3, we can estimate an upper bound on the variance reduction PEMC can achieve relative to standard MC under these idealized assumptions.
Lemma 4.
Assume $c_{f-g}=c_{f}$. In the same setup as Lemma 3, the variance ratio between PEMC under the optimal allocation and standard MC follows as
@@
Proof.
Proof Based on the results of Lemma 3, we can deduce that this ratio again does not depend on the budget $B$, so we omit its dependence in the statement. Moreover, we can conveniently choose a budget of the form $B=n_{0}c_{f-g}+n_{0}\frac{\sigma_{g}}{\sigma_{f-g}}\cdot\sqrt{\frac{c_{f-g}}{c_{g}}}c_{g}$ which gives an allocation of $n=n_{0}$ and $N=n_{0}\frac{\sigma_{g}}{\sigma_{f-g}}\cdot\sqrt{\frac{c_{f-g}}{c_{g}}}$ for PEMC in accordance with Lemma 3, while comparing with $B/c_{f}$ samples for standard MC. The rest follows from Lemma 2 and the assumption $c_{f}=c_{f-g}$. ∎
Based on the result of Lemma 4, a natural question arises: how can we gauge $\frac{\sigma_{f-g}^{2}}{\sigma_{f}^{2}}$ or $\frac{\sigma_{g}^{2}}{\sigma_{f}^{2}}$ in practice? How are these ratios tied to the quality of the predictive model $g$ in PEMC? As we shall see, the extent of variance reduction that PEMC can deliver is linked to how well $g$ is trained. To illustrate this, we first consider an ideal scenario where $g$ is trained to the optimum, i.e., the true regression function.
Lemma 5.
Suppose $c_{f-g}=c_{f}$, $f(\boldsymbol{Y})$ is square-integrable and $g=\mathbb{E}[f(\boldsymbol{Y})\,|\,\boldsymbol{X}]$. Define $\rho:=\mathrm{corr}(f,g)$ and $c:=\frac{c_{g}}{c_{f}}$. Then we have $\rho=\frac{\sigma_{g}}{\sigma_{f}}$ and the ratio $\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}$ in Lemma 4 reduces to
@@
Proof.
Proof When $g=\mathbb{E}[f(\boldsymbol{Y})\,|\,\boldsymbol{X}]$ is the true regressor, i.e.,
$g=\operatorname*{arg\,min}_{h\text{ measurable}}\mathbb{E}[(f(\boldsymbol{Y})-h(\boldsymbol{X}))^{2}],$
and $f$ is square-integrable, it follows that $\mathbb{E}[(f-g)h]=0$ for all square-integrable $h$. Plug in $h=g$, we obtain $\mathrm{Cov}(f-g,g)=0$. Consequently we have $\rho:=\mathrm{corr}(f,g)=\frac{\sigma_{g}}{\sigma_{f}}$ and $\sigma_{f}^{2}=\sigma_{f-g}^{2}+\sigma^{2}_{g}$, which further implies $\sigma_{g}=\rho^{2}\sigma_{f}^{2}$ and $\sigma_{f-g}^{2}=(1-\rho^{2})\sigma^{2}_{f}$. The rest follows from $c_{f-g}=c_{f}$. ∎
@@
Lemma 5 says that the variance reduction of PEMC relative to standard MC can be approximated by the function $r(\rho,c)$ in (10), where $\rho=\sigma_{g}/\sigma_{f}$ captures the predictive power of $g(\boldsymbol{X})$ for $f(\boldsymbol{Y})$, and $c=c_{g}/c_{f}$ quantifies the relative cost of evaluating $g(\boldsymbol{X})$ versus generating full samples for $f(\boldsymbol{Y})$. In Figure 1, we visualize the variance reduction function $r(\rho,c)$ for $\rho,c\in(0,1)$. On the left is a contour plot of $r(\rho,c)$ where the red line traces the curve $r(\rho,c)=1$, corresponding to the “break-even" boundary where PEMC’s variance matches that of MC. Regions below this line, $\{(\rho,r)\in[0,1]^{2}|r(\rho,c)<1\}$, represent regimes of $\rho$ and $c$ where PEMC achieves variance reduction, i.e., where $\rho$ is sufficiently large and $c$ is sufficiently small. On the right, the graph of $r(\rho,c)$ as a function of $\rho$ when $c=0.001$ provides a clear benchmark: a correlation $\rho=0.5$ yields approximately 22.2% variance reduction, while $\rho=0.7$ yields about 45.8%. In our PEMC applications, we design feature $\boldsymbol{X}$ and predictive model $g$ so that, for most $\boldsymbol{\theta}\in\Theta$ during evaluation, we can safely gauge that $c$ falls between $10^{-2}$ to $10^{-3}$ and $\rho$ exceeds 0.5. In these regions, Lemma 3 suggests choosing $N/n$ in the range of 5 to 20 for near-optimal performance. In our experiments, a ratio of $N=10n$ proved effective.
@@
Figure 1. Variance Reduction Function $r(\rho,c)$. Left: A contour map of the variance reduction ratio $r(\rho,c)$ as a function of the correlation $\rho=\sigma_{g}/\sigma_{f}$ and relative cost $c=c_{g}/c_{f}$. The red curve indicates the level set $r(\rho,c)=1$; the regimes where PEMC outperforms standard MC lie below the curve. Right: A graph of $r(\rho,c)$ as a function of $\rho$ when $c=10^{-3}$.
4.3.1. Learning Theory Estimates
@@
The statements in Lemma 3 assume that $g$ represents the true regression function $\mathbb{E}[f|\boldsymbol{X}]$. In practice, however, $g$ is obtained through a learning procedure applied to finite training samples, and thus will only approximate the true conditional expectation. Consequently, any practical implementation of PEMC must account for the approximation and estimation errors inherent in the ML training process, and this error also varies with $\boldsymbol{\theta}$. To rigorously quantify these effects, one can invoke tools from statistical learning theory, such as uniform convergence bounds, VC-dimension, or Rademacher complexities [126, 13, 94]. These quantities, by relating the complexity of the function class used to represent $g$ and the available training sample size $N_{\text{train}}$, lead to error bounds that ensure, with high probability, that the trained predictor $g$ is close to the true regression function within a controlled margin. We avoid delving deeply into them so as not to distract from our main theme. Instead, we adopt a standard, off-the-shelf and ‘placeholder’-type result, whose proof is detailed in the Appendix, that illustrates the typical form of guarantees one could expect from PEMC.
Lemma 6.
@@
4.4. Control Variate Coefficient
@@
$PEMC(a):=\frac{1}{n}\sum_{i=1}^{n}(f(\boldsymbol{Y}_{i})-ag(\boldsymbol{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}ag(\boldsymbol{\tilde{X}}_{j})$ (11)
@@
The introduction of $a$ is analogous to CV approaches, where the parameter $a$ does not introduce bias but can be tuned to minimize variance. In PEMC, this viewpoint may appear somewhat redundant since $ag$ is also a valid prediction model, which suggests $a$ would be chosen implicitly during the training of the prediction model. Nevertheless, suppose $g$ is fixed. Then $\frac{1}{n}\sum_{i=1}^{n}-g(\boldsymbol{X}_{i})+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\tilde{X}}_{j})$ can be viewed as a readily available, zero-mean control variate for $\frac{1}{n}\sum_{i=1}^{n}f(\boldsymbol{Y}_{i})$. We can then choose $a$ which minimizes the variance. The optimal $a^{*}$ then is:
$a^{*}=\frac{\mathrm{Cov}(f(\boldsymbol{Y}),g(\boldsymbol{X}))}{(n/N+1)\mathrm{Var}(g(\boldsymbol{X}))}.$
@@
In the ideal scenario, we have $N\gg n$ and $g(\boldsymbol{X})=\mathbb{E}[f(\boldsymbol{Y})|\boldsymbol{X}]$, which would imply $\mathrm{Cov}(f(\boldsymbol{Y})-g(\boldsymbol{X}),g(\boldsymbol{X}))=0$ and $\mathrm{Cov}(f(\boldsymbol{Y}),g(\boldsymbol{X}))=\mathrm{Var}(g(\boldsymbol{X}))$. This leads to $a^{*}=\frac{1}{1+n/N}\approx 1$, the default choice of PEMC, i.e., $a=1$.
5. Applications